OpenStudy (anonymous):

Let g be the function defined by g(x)=-x^2+2x g((squareroot)a),

OpenStudy (michele_laino):

try to substitute \[\sqrt{a}\] inplace of x

OpenStudy (michele_laino):

oops...in place of x

OpenStudy (michele_laino):

hint: \[g(\sqrt{a})=-[(\sqrt{a})^{2}]+2\sqrt{a}=...\] please continue

OpenStudy (anonymous):

YEah I got that part down but not sure how to solve after @Michele_Laino

OpenStudy (michele_laino):

please note that: \[(\sqrt{a})^{2}=a\]

OpenStudy (michele_laino):

so, you should get this: \[g(\sqrt{a})=-a+2\sqrt{a}\]

OpenStudy (anonymous):

Great thank you! @Michele_Laino I did not realize that (Sqareroot)a ^2 is = to a

OpenStudy (michele_laino):

thank you!